betty, carrie, and patty have the same birthday. Carry is now half of betty's age. Five years ago Patty was one-third of betty's age then. Which one will be 14 years old on her next birthday?
Answer by richwmiller(17219)

c=b/2
p-5=(b-5)/3
b=26 c=13 p=12
Carrie will be 14 years old on her next birthday.